OpenStudy (anonymous):

what is the solution of the system of equations? y= 3x + 7 y= x-9 Graph each system. tell whether the system has no solution, one solution, many solutions y= 2x - 3 y= -x +3

Parth (parthkohli):

We can say that 3x + 7 = x - 9 3x = x - 16 2x = -16 x = -8 Now, let's find y. y = x - 9 y = -8 - 9 y = -17

Parth (parthkohli):

(-8,-17)